A 0.300-mm-diameter silver wire carries a 10.0 mA current.

Part A

What is the electric field in the wire?

Express your answer with the appropriate units.

Part B

What is the electron drift speed in the wire?

Express your answer with the appropriate units.
Answer #1

Condictivity of silver = 6.3 × 107 =  

Also

So electric field = 1.764 x 10-3 V/m

Electron drift speed is found as

Free electron density of silver = 5.89 x 1028 m-3

So drift speed = I/(e n A) = 1.18 x 105 m/s

Please rate it up thanks :)